An object with an initial velocity of 20 m/s accelerates uniformly at 5 m/s 2 in the direction of its motion at a distance of 10 m. What is the final velocity of the object?



1
Expert's answer
2022-03-14T10:29:35-0400

"v_f^2=v_i^2+2as"


"v_f=\sqrt{v_i^2+2as}=\sqrt{(20\ m/s)^2+2\times5\ m/s^2\times10\ m}=22.4\ m/s"


Answer: 22.4 m/s